# Adaptive Designs in Fast-Track Registration Processes

### v1

@@ 5.3 Investigation of the second stage information for the case $Z_{1}\geq z_{f}$
Figure 8(b) compares the minimum second-stage information $t_{\xi}(I_{2,\min})$ (when $\xi=1.25$ and $\alpha_{c}=0.15$) for the case of a successful conditional registration ($Z_{1}\geq z_{f}$) between the designs. The minimum second-stage information is also represented in plots (a) to (d) of Figure 7 by the black solid lines. The blue dashed lines in these plots show for the case $Z_{1}\geq z_{f}$ the maximum second-stage information when it exceeds $I_{2,\min}$. The maximum second-stage information is given by the adaptive second-stage information $I_{2}^{A}(z_{f})$ (set $A\equiv\alpha$ for the non-adaptive design). These lines are linearly increasing in $t_{\xi}(I_{1})$ as long as $z_{f}=\Phi^{-1}(1-\alpha_{c})$, and they are constant or decreasing afterwards.
With a small interim information, it can happen that the maximum information $I_{2}^{A}(z_{f})$ becomes equal to the minimum information, in which case we need to use a constant second stage information also in the case $Z_{1}\geq z_{f}$, namely $I_{2,\min}$ (black solid line). For the non-adaptive design ($A\equiv\alpha$), this constant information equals $I_{\delta}$. Since $I_{2,\text{const}}=I_{\delta}$ also for $Z_{1}<z_{f}$, there is no difference between the lower and upper case in Figure 6. As a consequence, the pilot study is only done to "gamble" for a conditional registration. With $\xi=1.25$ and $\alpha_{c}=0.15$, this happens for all $t_{\xi}(I_{1})\leq 0.137$. Such a "gambling" with rather small pilot data appears to be a quite questionable strategy. The gambling area of the non-adaptive design is indicated in plot (a) of Figure 7 by the grey horizontal dashed line.
In our example, a constant second-stage information after the conditional registration occurs for all $t_{\xi}(I_{1})\leq 0.0813$ also with the inverse normal method, for all $t_{\xi}(I_{1})\leq 0.0854$ with Fisher’s product test, and for all $t_{\xi}(I_{1})\leq 0.1128$ with $A_{Z}$, i.e. for smaller pilot informations than in the non-adaptive case. For the adaptive designs the upper and lower cases can differ at least with regard to the constant second-stage information and the pilot data are always used in the application for permanent registration. Hence, the pilot data are never used only to "gamble" for a conditional registration.
@@ 5.4 Overall comparison with regard to the second stage information
Obviously, for the case $Z_{1}\geq z_{f}$, the data-driven second-stage information is always in the region between the solid black and dashed blue lines. One can see from Figure 7 that using a non-constant conditional error function requires much smaller second (and thereby overall) informations than the constant conditional error function $A\equiv\alpha$. With a sufficiently large pilot information, the region for the second-stage information can even be uniformly smaller for the adaptive designs than for the non-adaptive design. Moreover, also the constant information $I_{2,\text{const}}$ for case $Z_{1}<z_{f}$ becomes smaller with the non-constant conditional error functions. This clearly indicates the efficiency gain by using an adaptive design with non-constant conditional error function in comparison to a non-adaptive design with $A\equiv\alpha$.
Figure 9(a) allows for a comparison of the maximum information of the second stage between the designs. It can be seen that the non-adaptive design leads to a larger maximum information than the adaptive design with $A_{Z}$ if $t_{\xi}(I_{1})\geq 0.14$, with Fisher’s product test if $t_{\xi}(I_{1})\geq 0.17$ and with the inverse normal method if $t_{\xi}(I_{1})\geq 0.2$. A comparison of the maxmimum information over both stages can be found in the Appendix (Figure 17(a)).
Figure 9(b) shows for $\xi=1.25$ and $\alpha_{c}=0.15$ the average second stage information (relative to $I_{\delta}$) for the different designs when $\delta=\xi\delta_{\text{rel}}$. The non-adaptive design has a larger second-stage information than the inverse normal method for $t_{\xi}(I_{1})\geq 0.1398$ and Fisher’s product test for $t_{\xi}(I_{1})\geq 0.0742$. In particular, for the inverse normal method there is only a small non-gambling area for $t_{\xi}(I_{1})$ where the non-adaptive design has a smaller mean information than the inverse normal method. For Fisher’s product test and the adaptive design with $A_{Z}$ there is no such area. In particular, the adaptive design with $A_{Z}$ even has a lower mean information than the non-adaptive design for all considered $t_{\xi}(I_{1})$. A comparison of the maxmimum information over both stages can be found in the Appendix (Figure 17(b)).
@@
For our example fast-track procedure with assumption $\delta_{\text{rel}}=1.4$ and the considered relative first-stage information $t_{\xi}(I_{1})=0.5$, i.e. $n_{1}=0.5\cdot n_{\delta}\approx 69$, Table 1 shows the constant second-stage sample size per group $n_{2,\mathrm{const}}$ in case of no successful conditional registration and the minimum und maximum sample size per group in case of a successful conditional registration. The mean information over all $z_{1}\in\mathbb{R}$ and thus over both branches is also shown by $E(n_{2})$. The values in brackets show the quantities added by $n_{1}$, i.e. over both stages. Depending on the confidence of the applicant in the assumed a priori effect a comparison of the per group sample sizes of Table 1 with $n_{\delta}=137$ or $n_{\mathrm{rel}}:=2\hat{\sigma}^{2}\cdot\eta_{f}^{2}/\delta_{\text{rel}}^{2}=2\cdot 5.17^{2}\cdot 2.8^{2}/1.4^{2}\approx 214$ is interesting.
$A$ $n_{2,\mathrm{const}}$ $n_{2,\min}$ $n_{2,\max}$ $E(n_{2})$
@@
(a) Constant second stage information.
(b) Minimum second stage information
for upper branch.
Figure 8: Constant second stage information after non-successful conditional registration and minimum second stage information after successful conditional registration for $\xi=1.25$.
(a) Maximum information of second stage.
(b) Mean information of second stage.
Figure 9: Maximum and mean information for $\xi=1.25$ without assuming conditional registration as requirement for permanent registration.
@@ 5.5 Probability for a conditional registration
The probability for achieving conditional registration (i.e. $Z_{1}\geq z_{f}$) under the a priori assumed effect $\delta$ is shown by the green dot-dashed lines with values given by the right vertical axis. This curve is the same for all designs and plots of Figures 7, 8, 9. It illustrates the possibilities and limitations in reaching conditional registration for given first stage information $t_{\xi}(I_{1})$ when $\xi=\delta/\delta_{\text{rel}}$ is a small as $1.25$. For our example fast-track procedure the power for a conditional registration at $\delta=1.75$ is $\mathbb{P}_{\delta}(Z_{1}\geq z_{f})=0.65$.
@@ 5.6 Additional results and final remarks
In Appendix D, one can find the results of similar investigations for the alternative value of $\xi$, for example, for $\xi=\xi_{\min}$. The results are quite similar.
@@
We finally mention that there are several alternative strategies for designs with a data-driven decision to either apply for or waive the application for a conditional registration. For example, we could use an adaptive design with a benchmark boundary $z_{\text{s}}<z_{f}$, where for $z_{\text{s}}<Z_{1}<z_{f}$ one calculate the second stage information such that a conditional power of $1-\beta$ is achieved at the interim estimate $\max\{\delta_{\text{rel}},\hat{\vartheta}_{1}\}$, and to avoid a too large maximum second-stage information, a constant information is used for the second stage if $Z_{1}\leq z_{\text{s}}$, like in the lower branch of Figure 6. This information could then be calculated to meet the power $1-\beta$ conditional on $Z_{1}\leq z_{\text{s}}$. Furthermore, a binding futility bound less than $z_{f}$ could also be introduced, like a stopping if $\hat{\vartheta}_{1}\leq 0$, i.e. $Z_{1}\leq 0$, or stopping for futility, if a significant negative effect is observed in the pilot study. With such futility bounds we expect only small quantitative (and no qualitative) changes in the results and findings presented here, in particular, when the futility bound is non-binding.
Another, possibility is to introduce an early rejection boundary $\Phi^{-1}(1-\alpha_{1})$ for some $\alpha_{1}<\alpha$ for the first stage (pilot study) that, of course, would need to be accounted for in the choice of the conditional error function to still meet (9). This would provide the possibility to already achieve an permanent registration with the pilot study. However, we would think that in this case the pilot study would need to be sufficiently large to meet all requirements for a permanent registration.
@@
We finally would like to emphasise the importance of an overall type I error rate control when deciding in a data driven way whether to apply for conditional registration or waive this application and continue with a study for a permanent registration. Without such overall type I error rate control, we could end up with a substantially increased type I error rate, i.e. a probability for a permanent registration of an inefficient device that is much larger than $\alpha$. Assume that we start, like in Section 4, with a two-stage fast-track process using the binding futility rule $Z_{1}<z_{f}$ and a (non-constant) conditional error function $A$ at level $\alpha$. If, in the case of an unsuccessful attempt to register conditionally, we (always) start a new study for a direct, permanent registration at significance level $\alpha$, the overall type I error rate could become $(1+\Phi(z_{f}))\cdot\alpha$ which equals $0.04625$ for $\alpha=0.025$ and $z_{f}=\Phi^{-1}(1-\alpha_{c})$ with $\alpha_{c}=0.15$. We would conclude that this implies the requirement to disclose any attempt to continue after an unsuccessful pilot study with a study for a direct permanent registration. Moreover, if this intended, one would need to show that the anticipated strategy controls the overall type I error rate at the required level. An example for such a strategy was indicated in Figure 6 and discussed in this section.
@@ 6 Summary and Discussion
The main focus of this paper is on the utility of adaptive designs for a fast-track registration process like the DiGA fast-track process in Germany. This refers to the investigation of required minimum, maximum and average sample sizes under the aim of a certain overall success probability at an assumed a priori healthcare effect. Such an a priori assumption is not uncommon for adaptive designs, where we often speak of a sample size reassessment, indicating that some a priori sample size calculation has been performed (see, for instance, Chapter 7 in [10]). We have clearly seen that the implementation of the fast-track procedure with an adaptive design is more efficient than the widely adopted approach with separate studies. The use of adaptive designs often leads to smaller minimum, maximum and average sample sizes, because they permit to use the pilot data for the permanent registration study.
@@
Due to Section 4, adaptive designs are particularly more efficient than the independent study approach if an unsuccessful study for conditional registration precludes any further attempt to receive a permanent registration. In this case, adaptive designs are also less conservative than the separate study approach, because the latter has an effective overall type I error rate of $\alpha_{f}\cdot\alpha$ while the adaptive designs can achieve permanent registration with an overall type I error rate $\alpha$.
@@
We have seen in Section 3.2 that if an unsuccessful study for conditional registration precludes any further attempt to receive a permanent registration, a fast-track procedure is questionable when the a priori expectation on the health care effect does not go much beyond the minimal clinically relevant effect. This is because in these cases, the requirements for conditional registration are becoming more stringent than those for permanent registration. As a consequence, the sample size required for a successful conditional registration with a fast-track procedure then exceeds the sample size needed for a study that directly targets a permanent registration. As we have seen, this concerns all healthcare effects which are less than $43\%$ greater than the minimally clinically relevant effect if the one-sided level $\alpha=0.025$ is required and power $80\%$ is anticipated.
@@
This observation has motivated us to consider in Section 5 the second type of designs without a binding futility stop, which allow to decide based on the pilot data whether to apply for conditional registration or waive this application, and continue with a study for direct permanent registration. The presented designs are valid due to a strict control of the type I error rate. A prior announcement to the authorities of such a data-dependent mid-trial decision in advance ensures transparency. This appears to us to be an approach that is acceptable in a regulatory environment. Two different types of adaptive designs are presented in Section 5. The first type is based on a specific conditional error function selected for the fast-track, which is then extended by the possibility of skipping the conditional registration. The inverse normal method and Fisher’s product test are considered. The second type of design is based on the Müller and Schäfer method [27]. Starting with a test for direct permanent registration, the possibility of deciding data-dependent whether to apply for conditional registration is integrated. This is formalised by the conditional error function $A_{Z}$. Mathematically, the considered types of designs only differ in the way the conditional error function is selected. For the first type of design, the choice is ad hoc; for the other, the conditional error function of the permanent registration is used, which is in line with the Müller and Schäfer method. As we have seen, using the conditional error function $A_{Z}$ outperforms the non-adaptive design across all considered efficiency parameters for at least sufficiently large pilot studies and shows clear advantages when compared with the other adaptive designs.
An additional potential advantage of adaptive designs that has not been addressed in this paper is the robustness of the operation characteristics with regard to the a priori assumptions, the main reason to do a sample size recalculation at interim analyses or a pilot study. An investigation of the robustness of the adaptive designs in comparison to the separate stage approach would be an interesting topic for future investigations.
Other types of designs that also allow an early attempt to achieve permanent registration at the interim analysis are not considered here. This includes classical group sequential designs and adaptive designs with the possibility of early rejection of the null hypothesis while controlling the overall type I error rate (see, for instance, [10]). Such designs could be of particular interest when the sample sizes of the first stage are large, and so there might be sufficient evidence for a permanent registration; otherwise, the data could be used to apply for conditional registration. However, this requires the use of group sequential boundaries in order to account for the repeated attempts to reach permanent registration. Note that an adaptive design has the advantage of a classical group sequential design to permit a data-driven sample size calculation and the control of conditional power, which could be understood as a requirement for conditional registration.
At this point, it is important to emphasise that we do not consider a multiplicity adjustment with regard to conditional and permanent registration, i.e., we aim to control the probability of receiving either conditional or permanent registration (or both). We consider such an adjustment as unnecessary because of the temporary nature of the conditional registration that applies only to a rather short time interval (e.g. in Germany, one year, and in other countries, even less). Given this, it’s very likely that a patient will be affected by only one registration process, i.e. receives an inefficient health application due to either a temporary or permanent registration. This is in line with the idea of controlling only relevant type I error rates; see, for instance [29] and [30].
@@
This paper aims to illustrate the considerable potential of using adaptive designs in fast-track procedures such as those for digital health applications by presenting selected designs. However, this means that in certain situations, alternative designs may be more appropriate. For example, under the assumptions of Section 5, it is not necessary that both arms in Figure 6 achieve power $1-\beta$ for rejecting the null hypothesis conditional on the respective arm being realised. For instance, a conditional success probability $\pi>1-\beta$ could be assigned to the arm that covers successful conditional registration and a lower one to the other arm, such that an overall success probability of $1-\beta$ is still given. That could increase the success probability for conditional registration.
@@
Note that in Section 4, where conditional registration was a requirement for permanent registration, we can understand the case of a unsuccessful registration as a binding futility stop for the two-stage designs. Instead, we could have dealt with this case like an unbinding futility stop which would have led to a somewhat conservative design with somewhat larger sample sizes. However, such designs provide the opportunity to continue with a second stage for a permanent registration whenever there are good reasons for and sufficient resources have been achieved. This opportunity may outweigh the rather limited increase in the sample sizes. Corresponding graphs showing the efficiency of such designs can be found in appendix C.2.
@@
One should note that in most parts of the paper, the choice of the considered conditional error functions is somewhat arbitrary. Thus, there could be functions that lead to greater efficiency. For instance, one could consider the optimal conditional error functions suggested by Brannath and Bauer [31], which minimise the expected sample size under the sample-size recalculation approach to achieve a certain conditional power used in this paper.
@@
Finally, we would like to mention some important aspects of a fast-track registration process that remained unaddressed in this paper because it’s either out of the scope of our consideration or would have overloaded our investigation. Most of our discussion relies on the existence of a well-accepted minimally clinically relevant healthcare effect. As indicated in our running example this may not be the case in practice for all endpoints. Moreover, there are quite a few different approaches to determine a minimally clinically relevant effect, and there seems to be no clear guidance on which approach to use; see, for instance, [32], [33], [34]. Moreover, as we have seen, the requirement that the point estimator of the healthcare effect exceeds the minimally clinically relevant effect, can be stronger than the standard requirement of a p-value below $\alpha=0.025$ for sample sizes that are just as large as required for a power of $80\%$. This questions the use of a minimally clinically relevant effect. However, this does not apply to pilot or interim data as long as their sample sizes are moderate and hence a minimally clinically relevant effect may well be used as requirement for conditional registration, but probably not for the permanent registration. In any case, for our exposition, the assumption of a minimal clinically relevant effect provided a delightful context for a comparison of different designs for a fast-track procedure.
@@
Another unaddressed possibility is the use of surrogate endpoints for conditional registration, as it has been suggested in the pharmaceutical context; see, for instance, [35], [36]. This requires more discussions and investigations (see, for example, [37] and [38]), in particular when applied to digital health applications. Moreover, there is the possibility of selecting an endpoint at interim for conditional and permanent registration. Further investigations could also be conducted in this context.
@@
A further unaddressed important issue that goes completely beyond our discussion are health economic aspects that consider the balance between gains and costs of a fast-track procedure.

### v3

@@ 5.3 Investigation of the second stage information for the case $Z_{1}\geq z_{f}$
Figure 7 shows the minimum second-stage information $t_{\xi}(I_{2,\min})$ by the solid lines and the maximum second-stage information $t_{\xi}(I_{2,\max})$ by the dashed lines for the different designs for the case of a successful conditional registration ($Z_{1}\geq z_{f}$). The maximum second-stage information is given by the adaptive second-stage information $I_{2}^{A}(z_{f})$ (set $A\equiv\alpha$ for the non-adaptive design) when it exceeds $I_{2,\min}$. These lines are linearly increasing in $t_{\xi}(I_{1})$ as long as $z_{f}=\Phi^{-1}(1-\alpha_{c})$, and they are constant or decreasing afterwards.
With a small interim information, it can happen that the maximum information $I_{2}^{A}(z_{f})$ becomes equal to the minimum information, in which case we need to use a constant second stage information also in the case $Z_{1}\geq z_{f}$, namely $I_{2,\min}$ (solid line). For the non-adaptive design ($A\equiv\alpha$), this constant information equals $I_{\delta}$. Since $I_{2,\text{const}}=I_{\delta}$ also for $Z_{1}<z_{f}$, there is no difference between the lower and upper case in Figure 6. As a consequence, the pilot study is only done to “gamble” for a conditional registration. With $\xi=1.25$ and $\alpha_{c}=0.15$, this happens for all $t_{\xi}(I_{1})\leq 0.137$. Such a “gambling” with rather small pilot data appears to be a quite questionable strategy. The gambling area is indicated in plot (a) of Figure 7 by the vertical solid line. For the considered adaptive designs, a constant second-stage information after successful conditional registration occurs for smaller pilot informations than in the non-adaptive case. For the adaptive designs, the upper and lower cases can differ at least with regard to the constant second-stage information, and the pilot data are always used in the application for permanent registration. Hence, the pilot data are never used only to “gamble” for a conditional registration.
@@ 5.4 Overall comparison with regard to the second stage information
Obviously, for the case $Z_{1}\geq z_{f}$, the data-driven second-stage information is always in the region between the solid and dashed lines in Figure 7. One can see from this figure that using a non-constant conditional error function requires much smaller second (and thereby overall) informations than the constant conditional error function $A\equiv\alpha$. With a sufficiently large pilot information, the region for the second-stage information can even be uniformly smaller for the adaptive designs than for the non-adaptive design. Moreover, also the constant information $I_{2,\text{const}}$ for the case $Z_{1}<z_{f}$ becomes smaller with the non-constant conditional error functions. This clearly indicates the efficiency gain by using such adaptive designs.
Figure 8 (left) allows for a comparison of the maximum information of the second stage between the different designs over both branches. It can be seen that the non-adaptive design leads to a larger maximum information than the adaptive design with $A_{Z}$ if $t_{\xi}(I_{1})\geq 0.14$, with Fisher’s product test if $t_{\xi}(I_{1})\geq 0.17$ and with the inverse normal method if $t_{\xi}(I_{1})\geq 0.2$.
Figure 8 (right) shows for $\xi=1.25$ and $\alpha_{c}=0.15$ the average second stage information (relative to $I_{\delta}$) for the different designs when $\delta=\xi\delta_{\text{rel}}$. The non-adaptive design has a larger second-stage information than the inverse normal method for $t_{\xi}(I_{1})\geq 0.1398$ and Fisher’s product test for $t_{\xi}(I_{1})\geq 0.0742$. In particular, there is only a small non-gambling area for $t_{\xi}(I_{1})$ where the non-adaptive design has a smaller mean information than the inverse normal method. For Fisher’s product test and $A_{Z}$, there is no such area. The adaptive design with $A_{Z}$ even has a lower mean information than the non-adaptive design for all considered $t_{\xi}(I_{1})$. A comparison of the maximum and mean information over both stages can be found in the Supplement.
@@
For our example fast-track procedure with assumption $\delta_{\text{rel}}=1.4$ and the considered relative first-stage information $t_{\xi}(I_{1})=0.5$, i.e. $n_{1}=0.5\cdot n_{\delta}\approx 69$, Table 1 shows the constant second-stage sample size per group $n_{2,\mathrm{const}}$ in case of no successful conditional registration and the minimum and maximum sample size per group in case of a successful conditional registration. The mean information over all $z_{1}\in\mathbb{R}$ and thus over both branches is also shown by $E(n_{2})$. The values in brackets show the quantities added by $n_{1}$, i.e. over both stages. Depending on the confidence of the applicant in the assumed a priori effect, a comparison of the per group sample sizes of Table 1 with $n_{\delta}=137$ or $n_{\mathrm{rel}}:=2\hat{\sigma}^{2}\cdot\eta_{f}^{2}/\delta_{\text{rel}}^{2}=2\cdot 5.17^{2}\cdot 2.8^{2}/1.4^{2}\approx 214$ may be of interest.
AA $n_{2,\mathrm{const}}$ $n_{2,\min}$ $n_{2,\max}$ $E(n_{2})$
@@
Figure 8: Maximum information (left) and mean information (right) for $\xi=1.25$ over both branches.
@@ 5.5 Probability for a conditional registration
The probability for achieving conditional registration (i.e. $Z_{1}\geq z_{f}$) under the a priori assumed effect $\delta$ is shown by the dot-dashed lines with values given by the right vertical axis. This curve is the same for all designs and plots of Figures 7 and 8. It illustrates the possibilities and limitations in reaching conditional registration for given first-stage information $t_{\xi}(I_{1})$ when $\xi=\delta/\delta_{\text{rel}}$ is a small as $1.25$. For our example fast-track procedure, the power for a conditional registration at $\delta=1.75$ is $\mathbb{P}_{\delta}(Z_{1}\geq z_{f})=0.65$.
@@ 5.6 Additional results and final remarks
In the Supplement, one can find similar results of analogous investigations for the alternative value of $\xi=\xi_{\min}$.
@@
We finally mention that there are several alternative strategies for designs with a data-driven decision to either apply for or abstain from the application for a conditional registration. For example, we could use an adaptive design with a benchmark boundary $z_{\text{s}}<z_{f}$, where for $z_{\text{s}}<Z_{1}<z_{f}$ one calculates the second stage information such that a conditional power of $1-\beta$ is achieved at the interim estimate $\max\{\delta_{\text{rel}},\hat{\vartheta}_{1}\}$, and to avoid a too large maximum second-stage information, a constant information is used for the second stage if $Z_{1}\leq z_{\text{s}}$, like in the lower branch of Figure 6. This information could then be calculated to meet the power $1-\beta$ conditional on $Z_{1}\leq z_{\text{s}}$. Furthermore, a binding futility bound less than $z_{f}$ could also be introduced, like a stopping if $\hat{\vartheta}_{1}\leq 0$, i.e. $Z_{1}\leq 0$, or stopping for futility, if a significant negative effect is observed in the pilot study. With such futility bounds, we expect only small quantitative (and no qualitative) changes in the results and findings presented here, in particular, when the futility bound is non-binding.
@@
Moreover, we would like to emphasize the importance of an overall type I error rate control when deciding in a data-driven way whether to apply for conditional registration or abstain from this application and continue with a study for permanent registration. Without such overall type I error rate control, we could end up with a substantially increased probability for a permanent registration of an inefficient device that is much larger than $\alpha$. Assume that we start, like in Section 4, with a two-stage fast-track process using the binding futility rule $Z_{1}<z_{f}$ and a (non-constant) conditional error function $A$ at level $\alpha$. If, in the case of an unsuccessful attempt to register conditionally, we (always) start a new study for a direct, permanent registration at significance level $\alpha$, the overall type I error rate could become $(1+\Phi(z_{f}))\cdot\alpha$ which equals $0.04625$ for $\alpha=0.025$ and $z_{f}=\Phi^{-1}(1-\alpha_{c})$ with $\alpha_{c}=0.15$. We would conclude that this implies the requirement to disclose in advance the attempt to continue after an unsuccessful pilot study with a study for direct permanent registration. Moreover, if this is intended, one would need to show that the anticipated strategy controls the overall type I error rate at the required level. An example of such a strategy was indicated in Figure 6 and discussed in this section.
@@ 6 Summary and Discussion
The main focus of this paper is on the utility of adaptive designs for fast-track registration processes (for digital health applications) like the DiGA fast-track in Germany. This refers to the investigation of required minimum, maximum, and average sample sizes under the aim of a certain overall success probability at an assumed a priori healthcare effect. Such an a priori assumption is not uncommon for adaptive designs, where we often speak of a sample size reassessment, indicating that some a priori sample size calculation has been performed (see, for instance, Chapter 7 in the work of Wassmer and Brannath [15]). We have clearly seen that the implementation of the fast-track procedure with an adaptive design is more efficient than the adopted approach with separate studies, whereby we worked with the assumption of the same design, same primary endpoints, and same study population for conditional and final registration. The use of adaptive designs often leads to smaller minimum, maximum, and average sample sizes, because they permit the use of the pilot data for the permanent registration study.
@@
Due to Section 4, adaptive designs are particularly more efficient than the independent study approach if an unsuccessful study for conditional registration precludes any further attempt to receive a permanent registration. In this case, they are also less conservative than the separate study approach that has an effective overall type I error rate of $\alpha_{f}\cdot\alpha$ while the adaptive designs can achieve permanent registration with an overall type I error rate $\alpha$.
@@
We have seen in Section 3.2 that if an unsuccessful study for conditional registration precludes any further attempt to receive a permanent registration, a fast-track procedure is questionable when the a priori expectation on the health care effect does not go much beyond the minimal clinically relevant effect. This is because in these cases, the requirements for conditional registration are becoming more stringent than those for permanent registration.
@@
This observation has motivated us to consider in Section 5 the second type of designs without a binding futility stop, which allow to decide based on the pilot data whether to apply for conditional registration or abstain from this application, and continue with a study for direct permanent registration. The presented designs can be considered for all a priori effect assumptions and are valid due to a strict control of the type I error rate. A prior announcement to the authorities of such a data-dependent mid-trial decision in advance ensures transparency. This appears to us to be an approach that is acceptable in a regulatory environment. Two different types of adaptive designs are presented in Section 5. The first type is based on a specific conditional error function selected for the fast-track route, which is then extended by the possibility of skipping the conditional registration. The inverse normal method and Fisher’s product test are considered. The second type of design is based on the approach of Müller and Schäfer [33]. Starting with a test for direct permanent registration, the possibility of deciding data-dependent whether to apply for conditional registration is integrated. This is formalized by the conditional error function $A_{Z}$. Mathematically, the considered types of designs only differ in the way the conditional error function is selected. For the first type of design, the choice is ad hoc; for the other, the conditional error function of the permanent registration is used, which is in line with the Müller and Schäfer method. As we have seen, using the conditional error function $A_{Z}$ is convincing and outperforms the non-adaptive design across all considered efficiency parameters for at least sufficiently large pilot studies.
@@
As already mentioned, this paper aims to illustrate the considerable potential of using adaptive designs in fast-track procedures such as those for digital health applications by presenting selected designs. However, this means that in certain situations, alternative designs may be more appropriate. For example, under the assumptions of Section 5, it is not necessary that both arms in Figure 6 achieve power $1-\beta$ for rejecting the null hypothesis conditional on the respective arm being realized. For instance, a conditional success probability $\pi>1-\beta$ could be assigned to the arm that covers successful conditional registration and a lower one to the other arm, such that an overall success probability of $1-\beta$ is still given. That could increase the success probability for conditional registration.
@@
Note that in Section 4, where conditional registration was a requirement for permanent registration, we can understand the case of an unsuccessful registration as a binding futility stop for the two-stage designs. Instead, we could have dealt with this case like an unbinding futility stop, which would have led to a somewhat conservative design with somewhat larger sample sizes. However, such designs provide the opportunity to continue with a second stage for permanent registration whenever there are good reasons for and sufficient resources have been achieved. This opportunity may outweigh the rather limited increase in the sample sizes. Corresponding graphs showing the efficiency of such designs can be found in the Supplement.
@@
One should note that in most parts of the paper, the choice of the considered conditional error functions is somewhat arbitrary. Thus, there could be functions that lead to greater efficiency. For instance, one could consider the optimal conditional error functions suggested by Brannath and Bauer [35], which minimize the expected sample size under the sample-size recalculation approach to achieve a certain conditional power used in this paper.
An additional potential advantage of the adaptive designs mentioned in this paper is the robustness of the operation characteristics with regard to the a priori assumptions, the main reason to do a sample size recalculation at interim analyses or a pilot study. An investigation of the robustness of the adaptive designs in comparison to separate stage approaches would be an interesting topic for future investigations.
Other types of designs that also allow an early attempt to achieve permanent registration at the interim analysis are not considered here. This includes classical group sequential designs and adaptive designs with the possibility of early rejection of the null hypothesis while controlling the overall type I error rate (see, for instance, the work of Wassmer and Brannath [15]). Such designs could be of particular interest when the sample sizes of the first stage are large, and so there might be sufficient evidence for a permanent registration; otherwise, the data could be used to apply for conditional registration. However, this requires the use of group sequential boundaries in order to account for the repeated attempts to reach permanent registration. Note that an adaptive design has the advantage over a classical group sequential design to permit a data-driven sample size calculation and the control of conditional power, which can be understood as a requirement for conditional registration.
At this point, it is important to emphasize that we do not consider a multiplicity adjustment with regard to conditional and permanent registration, i.e., we do not aim to control the probability of erroneously receiving either conditional or permanent registration (or both). We consider such an adjustment as unnecessary because of the temporary nature of the conditional registration that applies only to a rather short time interval (e.g. in Germany, one year, and in other countries, even less). Given this, it’s very likely that a patient will be affected by only one registration process, i.e. receives an inefficient health application due to either a temporary or permanent registration. This is in line with the idea of controlling only relevant type I error rates; see, for instance the works of Brannath et al [36] and Brannath [37].
@@
Finally, we would like to mention some important aspects of a fast-track registration process that remained unaddressed in this paper because it’s either out of the scope of our consideration or would have overloaded our investigation. Most of our discussion relies on the existence of a well-accepted minimally clinically relevant healthcare effect. As indicated in our running example, this may not be the case in practice for all endpoints. Moreover, there are quite a few different approaches to determine a minimally clinically relevant effect, and there seems to be no clear guidance on which approach to use; see, for instance, the works of Devji et al [38], Franceschini et al [39] and Vach and Saxer [40]. Moreover, as we have seen, the requirement that the point estimator of the healthcare effect exceeds the minimally clinically relevant effect can be stronger than the standard requirement of a p-value below $\alpha=0.025$ for sample sizes that are just as large as required for a power of $80\%$. This questions the use of a minimally clinically relevant effect as a requirement for permanent registration. However, this does not apply to pilot or interim data as long as their sample sizes are moderate, and hence a minimally clinically relevant effect may well be used as a requirement for conditional registration. In any case, for our exposition, the assumption of a minimal clinically relevant effect provided an insightful context for a comparison of different designs for a fast-track procedure.
@@
Another unaddressed possibility is the use of surrogate endpoints for conditional registration, as it has been suggested in the pharmaceutical context [41, 42]. This requires more discussions and investigations (see, for example, the works of Shih et al [43] and Shih [44]), in particular when applied to digital health applications. Moreover, there is the possibility of selecting an endpoint at interim for conditional and permanent registration. In particular, the endpoints as well as the study populations may differ between conditional and permanent registration. Further investigations could also be conducted in this context.
@@
A further unaddressed important issue that goes completely beyond our discussion of our paper are health economic aspects that consider the balance between gains and costs of a fast-track procedure.
Acknowledgments
The authors kindly thank PD Dr. Norbert Benda for helpful discussions and comments. Further, the authors gratefully acknowledge the support of the Leibniz ScienceCampus Bremen Digital Public Health (www.digital-public-health.de), which is jointly funded by the Leibniz Association (W72/2022), the Federal State of Bremen, and the Leibniz Institute for Prevention Research and Epidemiology – BIPS.